A Complete Bibliography of Publications in Journal of Computational Physics: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.03

Title word cross-reference

2 [BR69]. 3 [BR69]. K_{p}(x) [Aha69]. S [Car69a].

-State [Car69a].

accuracy [HS68b, Lat69]. acetic [MB69]. acid [MB69]. acoustic [HJ69, ZD67]. adapted [HS68a]. advection [Cro67b]. aided [HS68a]. algebra [Cam68, Lev67, Ste67, Swa69]. Algol [Ber68]. algorithms [Swa69]. almost [Cro67a, HA68]. alphabetic [GK69]. among [KKS68]. analysis [Cam66, Gra68, Kop69, TW69]. analytic [Sch69a]. analytical [JK69]. analyze [SG69]. angle [WCT68]. angles [Yut67]. anisotropic [vdH68]. appearing [Che69]. applicable [Bye67]. Application [BD69, Les68, LC68, NH69]. approach [BG69b]. Approximation [FJ68, Dor68, Pie67]. approximations [BK67, Wil66]. arbitrarily [Tay68]. arbitrary [Ste68, Vie69]. argument [Zak69]. arising [Les68, LC68, SR68]. artificial [Cam66, vL69]. associated [Ste67]. astron [KR66]. Asymptotic [BG69b, vdH68]. atmosphere [Pie67, Yan69]. atmospheres [GB66]. atmospheric [SR68]. atom [HM69]. atomic [JK69, Kal66, SH68]. Author [Ano67a, Ano68, Ano69a, Ano69b]. automation [Gla68].

bases [Wol68]. basic [KK67]. behavior [ZD67]. Berkeley [GLSM68]. beta [Bye67]. body [BD69, BF69]. Boltzmann [HS68b, LC66]. Born [Sch66]. both [BK67]. bound [KE68]. bound-state [KE68]. boundaries [Vie69]. boundary [AWS69, KK67, LC68, Mat68, RR68]. boundary-layer [LC68]. boundary-value [KK67, Mat68, RR68]. bubble [Yut67]. bubble-chamber [Yut67].

calculate [Way68]. calculating [AI69, DS67, Tro66a, Tro66b, WCT68]. Calculation [Zak69, Cro67a, Gab68, HA68, Lap67, Mat69, Tay68]. calculations [BR69, Bur66, Car69b, Dal69, Del68, HM69, HS68c, Sau67]. calibrating [Ber68]. Carlo [BKKV69, HS68b, LJS69a, LJS69b, Mat69, TW69]. caused [Cam66]. cells [BF69]. certain [AWS69, Bak69, Wil66]. chains [Ver69]. chamber [Yut67]. characteristics [Sau67]. charged [Tro66b]. chemical [GLSM68, Tro66b]. chemistry [AF68, Ham68]. chemists [Ham68]. choice [Rig69]. circular [WCT68]. circulation [Bry69]. class [AWS69, BK67]. closed [FSD^{+}67]. closed-loop [FSD^{+}67]. Clouds [BF69]. clouds-in-cells [BF69]. Clouds-in-clouds [BF69]. coalescence [AJ69]. coefficients [NH69]. coil [Ver69]. collection [GLSM68]. collisions [Way68]. combined [CS67]. Comment [Lau69]. Comments [Yee70]. compact [Har69]. comparison [Gar69]. complementary [Zak69]. complex [KE68, Zak69]. compressible [FA68, GMD66, RB67, vL69]. Computation [Aha69, Hag68, Kop69, BT69, Che69, Har69, KR66, Ste68]. Computational [Ara66, Ste67, BK67]. Computer [AF68, HS68a, Swa69, Bye67, FSD^{+}67, HH69, Lev67, PR67]. Computer-aided [HS68a]. Computers [Ham68]. computing [AJ69, KM69b, Vie69]. concerning [Bla67]. condition [RR68]. conditional [FJ68]. conditions [HS68c, Tay69]. conducting [Kel67]. conduction [CS67]. conductivity [JPW68]. conference [Ham68]. confined [TLF69]. Conservation [APR68]. Construction [SH68, HS68a]. containing [Bur66, Tay68]. content [Sch66]. control [AF68, GLSM68]. convective [Fro68]. convergence [RR68]. Coordinate [Ush67, Ush68, APR68]. coordinates [BT69, LJS69a]. corrective [HH67]. correlation [JK69]. coupled [Hat69, Les68]. coupling [PR67]. cross [Ste68]. crystallography [GK69, Ham68]. crystals [Kop69]. cubic [Gab68]. curve [Che69]. curved [Mat68]. curvilinear [APR68]. cylindrical [Kel67].

data [GLSM68, Sou69, Win68]. decays [Tro66a]. degenerate [LC69]. derivatives [Les68]. description [RR68]. design [Ara66, BS68]. detached [Lap67]. Determination [BG69a, Yut67, MB69]. diatomic [SG69].

Difference [Deb67, PH68, RB67, BK67, Bun67, Bur66, Dor68, Fro68, Gar69, HP68, Hir68, KL69, Wen69, Wil66, vL69]. differences [Lap67]. differencing [Eme68, GMD66, Lat69]. differential [Hag68, KL69, Kus69, Les68, SR68, Slo68c, Wil66]. diffuse [GB66]. diffusion [vL69]. digital [BS68, FSD^{+}67, Lev67, PR67]. dimensional [AI69, APS66, Ara66, BT69, Cos69, HP68, JPW68, O’B69, PH68, Yee70]. dimensions [BR69, Bye67, Con67]. dimers [MB69]. dipole [DS67]. Dirac [Cam68, Lev67, Swa69]. direct [Ng68, O’B69, Yee70]. Dirichlet [Lic69, Tay69]. discontinuities [Bur66]. discontinuous [KH69]. Discrete [BD69, FA68, Gra68]. disk [WCT68]. disks [HH69]. dispersion [Fro68]. distribution [AJ69]. distributions [BKKU67]. divergence [Lic69]. domain [BS68]. due [AJ69]. dynamic [Tro66a]. Dynamics [ZD67, AH68, KH69, Wil66].

Effect [RR68]. effects [Dal69]. eigenfunctions [SH68]. eigenvalue [LC68]. eigenvalues [Hag68, KE68]. elastic [PH68]. electric [Ste68]. electrode [Kel67]. electron [HM69, KR66]. electron-atom [HM69]. Electronic [CD66]. elements [Car69a, HS68a, PR67, Tro66b]. elliptic [Mat68]. energy [FSD^{+}67, MB69, Tro66b, Way68]. entrainment [Cro67a]. equation [BD69, Bla67, Con67, Deb67, Dor68, GS67a, GS67b, GS69, KK67, KM69a, Lat69, Les68, O’B69, RO69, Win66, Yee70, HS68b, LC66]. equations [APR68, Ara66, CS67, Deb67, Hag68, Hat69, Hir68, KF68, KE68, KL69, Kus69, Les68, MR68, PH68, RB67, SR68, Slo68c, Wil66, vL69]. equilibrium [Ush67]. equivalences [KKS68]. Erratum [Ano67b]. Error [GB68, Zak69]. errors [Cam66, Del68, Slo68a]. estimates [GB68]. estimation [BKKU67]. Eulerian [FA68, GMD66]. evaluating [NH69]. Evaluation [LN69, Eme68, GCO67, JK69, LJS69a, PR67, Slo68b]. evaporation [AJ69]. events [BKKV69]. example [Inn66, KH69]. exchange [LN69]. excitations [ZD67]. expansion [Ush68]. experimental [Mil67, Sou69]. experiments [GLSM68]. explicit [Les68]. exterior [Dor68]. external [BKKU67, Vie69].

Feynman [LJS69a, LJS69b]. field [BKKU67, Con67, Tay69, Win68, Yut67]. fields [DS67, Ste68]. filters [BS68]. Finite [Bur66, Dor68, BK67, BS68, Hir68, Lap67]. Finite-difference [Bur66, Dor68, BK67, Hir68]. finite-memory [BS68]. first [Les68]. Fitting [Sou69, vdH68]. flow [Ara66, BT69, Cho67, Eme68, GMD66, HA68, HS68c, NH69, PH68, TW69, vL69]. flows [BG69b, Bur66, Cro67a, Sau67, Tay68]. fluid [AH68, APS66, Ara66, Eme68, KH69, Vie69, Wil66]. Flux [KKS68]. Fock [MR68]. Fokker [KF68]. form [APR68, BK67]. formation [KR66]. formed [KF68]. Formic [MB69]. FORTRAN [Tro66a, Tro66b, AJ69, DS67, SG69, WCT68, Wol68]. Fourier [GK69, GCO67]. Free [HS68c]. Free-surface [HS68c]. frontal [AI69]. function [Inn66, Kal66, Zak69]. functions [Car69b, HS68a, Ng68, Sch69a]. gap [RB69]. gas [BT69, NH69, RB67, Tay68]. gasdynamic [Cro67a]. general [BK67, DS67, HH67]. generating [BKKV69]. global [Cro68]. glow [Che69]. gravity [HJ69]. group [Wol68].

Hamiltonian [KM69b, PR67]. handling [Wol68]. harmonics [Gab68]. Hartmann [BG69b]. Hartree [MR68]. heat [CS67]. heat-transfer [CS67]. helium [Kal66]. Helmholtz [Yee70, O’B69]. Heuristic [Hir68]. high [FSD^{+}67, PR67, Way68]. high-energy [FSD^{+}67, Way68]. high-resolution [PR67]. higher [Ng68]. hydrodynamic [Bur66, Dal69]. hydrodynamics [APR68]. hydrogen [KF68]. hyperbolic [HP68].

II [GS67b, Ush68]. imaginary [GS67a, GS67b, LJS69a, LJS69b]. imbedding [BKK66, Hag68]. including [Dal69, Vie69]. incompressible [Ara66, Cho67, HA68, HS68c, Vie69]. incompressible-flow [HS68c]. incorporating [BK67]. index [Ano67a, Ano68, Ano69a, Ano69b]. infinity [Tay69]. Information [Sch66]. infrasonic [Pie67]. Initial [KK67, HH67, KL69, Kus69, Mil67]. Initial-value [KK67, HH67, KL69, Kus69, Mil67]. injection [KF68]. integral [Che69, FJ68, KK67, KM69a]. Integrals [HM69, GCO67, JK69, LN69, Sch69b, Slo68b, LJS69a]. Integration [GS67a, GS67b, Ara66, LJS69b, Les68, Sch69a]. interaction [MB69]. interactions [Tro66a]. interchange [Bye67]. interface [Ush67, Ush68]. internal [DS67]. interpolation [Inn66]. Invariant [BKK66]. investigations [APS66]. inviscid [Eme68, RB67, vL69]. involving [Ng68]. isothermal [Yan69]. isotropic [KKS68]. iteration [Mat68]. iterative [GB68].

kinematic [Tro66a]. Knudsen [TW69]. Kutta [Slo68a].

L [SH68]. L-S [SH68]. laboratory [Gla68, AF68]. Lagrangian [Cro67a]. large [CD66, Tay68]. lattices [DS67, JPW68, ZD67]. Lawrence [AF68]. layer [KR66, LC68]. layers [vdH68]. least [BD69]. Legendre [Deb67]. line [FSD^{+}67, GLSM68, Kop69]. linear [Slo68c]. linearization [LC68]. linearized [HJ69, NH69]. Localized [ZD67]. location [Ush67, Ush68]. long [Ara66]. long-term [Ara66]. loop [FSD^{+}67]. low [Ber68, Bye67]. low-beta [Bye67]. Luke [NH69].

MAC [Vie69]. magnetic [Ste68, Win68, Yut67]. magnetostatic [Con67]. magnetostatic-field [Con67]. many [BF69]. many-body [BF69]. maps [GK69]. mass [Cro67a]. matrix [Cam68, Car69a, HS68a, Lev67, PR67, Sch67]. mean [SR68]. measurements [BKKU67]. mechanics [Sch67]. media [Cos69, FA68]. medium [Kel67]. memory [BS68]. mesh [Win66]. Method [Slo68c, AI69, BD69, Bry69, BKKV69, Cho67, Fro68, GMD66, GK69, GS69, Har69, Les68, Mil67, SR68, Sau67, Vie69, vdH68]. Methods [GCO67, Eme68, Gra68, KK67, KL69, Kus69, Rig69, RB67, Slo68a]. mixed [Mat68]. Model [Mat69, APS66, Bye67, Cro68, FA68, GB66, HH69, HJ69, Ver69, Way68]. models [KKS68, Ush67]. modes [Bye67]. modifications [AJ69]. molecular [CD66]. molecule [Car69b]. molecules [Car69a, SG69]. momentum [Yut67]. monitor [Gla68]. Monte [BKKV69, HS68b, LJS69a, LJS69b, Mat69, TW69]. motion [AI69, Ara66, Ver69]. multidimensional [Sau67]. multilayer [Pie67]. Multiple [MR68]. Muonic [Car69b, Car69a]. Muonic-molecule [Car69b].

natural [BT69]. neutral [KF68]. NMR [Kop69, PR67]. Non [LN69]. Non-exchange [LN69]. nonlinear [AWS69, BK67, Bye67, Con67, Cos69, HS68b, Lic69, ZD67]. nonlocal [RO69]. nonuniform [Win66, Yut67]. nonunique [KH69]. Note [Cam68, Slo68a]. nuclear [BD69]. number [Tay68]. Numerical [APS66, BKKU67, BT69, Con67, CS67, Gra68, HA68, KF68, LC66, Mat68, Sau67, Sch67, Sch69a, Sha67, Ste68, Win66, AH68, Ara66, Bry69, Cho67, Cro67b, Cro68, Gar69, GCO67, GS69, HH67, HJ69, KM69a, Kus69, Les68, Mil67, O’B69, RB69, Slo68c, Slo68b, Tay69, Yan69, Yee70]. Numerov [Slo68a].

ocean [Bry69, Cro68]. off [Del68]. On-line [GLSM68, FSD^{+}67]. one [APS66, HP68, JPW68, O’B69, WCT68, Yee70, AI69]. one-dimensional [APS66, HP68, JPW68, O’B69, Yee70]. optical [ZD67]. optimization [BR69]. order [Cro67b, Hag68, Lap67, Les68, SR68, Slo68c]. ordinary [SR68]. ordinate [Gra68]. other [KKS68].

paramagnetic [KM69b]. parametric [KM69b]. Part [Ara66, Cro68]. partial [BG69a, KL69, Kus69, Wil66]. partially [Kel67]. particle [BR69, Tro66a, Way68]. particles [Tro66b]. Path [LJS69a]. perfectly [Kel67]. peripheral [BKKV69]. perturbation [LC69]. phase [Mat69]. physics [BF69, FSD^{+}67]. pipe [Cro67a]. PL [Ush68]. Planck [KF68]. plane [Sha67, Tay68]. planetary [GB66]. plasma [BF69, Bye67, KF68]. plotting [GK69]. point [AWS69]. points [Bla67]. Poiseuille [BG69b]. Poisson [Dor68, Win66]. polar [LJS69a]. polymer [Ver69]. polynomial [Deb67]. Positivity [Lat69]. potential [Kel67, RO69]. Practical [Bla67]. precise [Inn66]. prediction [Gar69]. principal [Slo68b]. principal-value [Slo68b]. principles [Sch69b]. problem [Bak69, BD69, Dor68, KK67, LC68, Lic69, Mat68, NH69, RR68]. problems [AWS69, Cho67, Eme68, GMD66, HH67, Mil67, Ste67, Tay69, TLF69, vdH68]. procedure [HH67]. procedures [Bun67]. Processing [Win68]. produced [Way68]. program [AJ69, Ber68, Cam68, DS67, KM69b, SG69, Tro66a, Way68, WCT68]. propagating [Yan69]. propagation [Cos69, Pie67]. propagator [LJS69b]. proton [Way68]. proton-proton [Way68]. pseudoviscosity [HP68]. PUFL [Cro67a]. pulsed [Kop69]. pulsed-NMR [Kop69].

quantities [Tro66a]. quasi [LC68, Ush67]. quasi-linearization [LC68]. quasi-static [Ush67]. quasilinear [Win66].

Racah [Ste67]. Radiation [AF68, BKKU67, CS67, GB66, ZD67]. radiation-field [BKKU67]. radiative [BKK66, KK67, Wen69]. raindrop [AJ69]. random [Rus68, Ver69]. random-coil [Ver69]. range [Tro66b]. range-energy [Tro66b]. rarefied [NH69]. ray [GK69]. Rayleigh [KKS68, LC69, NH69]. real [KE68]. realization [Rus68]. recursive [JK69]. reducing [Fro68]. reduction [GLSM68]. reflection [GB66]. reflections [Sha67]. region [Mat68]. relation [Tro66b]. relativistic [KR66]. relaxation [Rig69]. represent [Cam66]. representation [Deb67]. representations [Wol68]. required [LN69]. resistance [Ber68]. Resolution [Hat69, PR67]. resonance [KM69b]. results [BKKU67]. reversible [Bun67]. Riemann [Har69]. rotational [SG69]. Round [Del68]. Round-off [Del68]. Runge [Slo68a].

S [SH68]. scattering [HM69, KKS68, LN69]. scheme [Wen69]. schemes [Fro68, Gar69, HP68, vL69]. Schrödinger [BD69, Bla67, GS67a, GS67b, GS69, Hat69, LC69, Les68, RO69]. Second [Cro67b, Hag68, Lap67, Les68, SR68, Slo68c]. Second-order [Cro67b, Hag68, Lap67, Les68, SR68, Slo68c]. section [Ste68]. series [Ng68, Sch66]. set [Wol68]. several [Eme68]. shapes [Kop69]. sharing [Gla68]. shells [BKK66]. shock [Cam66, Lap67, Sha67]. shocks [Tay68]. short [Cam68]. simple [APS66, Gla68]. simulation [BF69, Ver69]. size [AJ69]. solid [WCT68]. Solution [AWS69, RO69, TLF69, BD69, Bla67, Con67, CS67, GS69, HH67, KM69a, KF68, KE68, LC66, O’B69, RR68, SR68, Sha67, Slo68c, Win66, Yee70]. solutions [BG69a, HS68b, KH69, MR68]. solving [Cho67, vdH68]. source [BKKU67]. Spatial [Lat69]. spectra [KM69b, Way68]. spectral [BK67]. spherical [BKK66, LJS69a]. spherically [FA68]. spin [KM69b, PR67]. spin-coupling [PR67]. spin-Hamiltonian [KM69b]. spline [Inn66]. squares [BD69]. Stability [KL69, Wil66, Hir68, Kus69, Mil67]. Stabilization [vL69]. Star [Ush67]. stars [HH69]. starting [Rig69]. State [Car69a, AJ69, Cam66, KE68]. static [Ush67]. statistical [Way68]. steady [AJ69, BT69, Cam66, Tay68]. steady-state [AJ69, Cam66]. Stochastic [Kal66]. stratified [Pie67]. stress [HS68c]. stretching [Ush67, Ush68]. strong [ZD67]. structure [CD66, SG69]. study [Bry69, RB69, Ver69, Yan69]. studying [KR66]. subdominant [BG69a]. submerged [Kel67]. subroutine [Tro66b]. subroutines [Wol68]. subtended [WCT68]. summation [Ng68]. supersonic [BT69, Tay68]. surface [Dal69, HS68c]. surfaces [MB69, Rus68]. symbolic [Har69]. symmetric [FA68]. symmetry [HS68a]. system [Gla68, Hat69, SR68]. systems [APR68, BK67, CD66, Kus69].

Taylor [RB69]. technique [Dal69, Tay69, Vie69]. techniques [Hag68, Sch67]. temperature [Pie67, Way68]. temperature- [Pie67]. temperatures [Ber68]. tension [Dal69]. tensor [Har69, Lev67]. term [Ara66]. terms [Cam66]. testing [Mil67]. tests [AI69]. theorem [Lic69]. theory [Che69, Hir68, LC68, LN69, SR68]. Thermal [JPW68]. thermometers [Ber68]. thick [vdH68]. three [BD69]. three-body [BD69]. Time [BS68, Bun67, Gla68, GS67a, GS67b, LJS69a, LJS69b]. Time-domain [BS68]. Time-reversible [Bun67]. time-sharing [Gla68]. tracks [Yut67]. transcendental [Ng68]. transfer [BKK66, CS67, KK67, Wen69, vdH68]. transient [AJ69]. transitions [Mat69]. Transport [AH68, LC66, Deb67, Lat69]. treating [Tay69]. triangle [Win66]. turbulence [AH68]. Two [Ara66, Cos69, AWS69, BT69, Bye67, Con67, Gar69, PH68, SR68, Way68]. Two-dimensional [Ara66, BT69, PH68]. two-point [AWS69]. two-temperature [Way68]. type [KL69, LN69].

undistorted [GK69]. unfolding [Bak69, GB68]. uniform [Ste68]. unsteady [GMD66, Sau67]. use [FSD^{+}67, SH68]. used [Gar69]. using [BG69a, Cam66, Hag68, PR67].

value [AWS69, HH67, KK67, KL69, Kus69, Mat68, Mil67, RR68, Slo68b]. Variable [HP68]. variation [Car69b]. Variational [Sch69b, Del68]. vectors [Rig69]. versatile [KM69b]. vertically [Yan69]. viscosity [Cam66]. viscous [Cho67, RB67, Sha67, Yan69]. Vogelaere [Les68]. Volterra [KM69a]. volume [Ano67a, Ano68, Ano69a, Ano69b]. vortex [TLF69]. vortices [RB69]. vs [Lat69].

wave [Cos69, HS68a, Kal66, KE68, Pie67]. waveguide [Ste68]. waves [Cam66, HJ69, SR68, Yan69]. weather [Gar69]. weight [NH69]. wide [RB69]. wide-gap [RB69]. Wiener [FJ68]. wind [Pie67]. wind-stratified [Pie67]. without [Les68]. world [Bry69]. Wronskian [BG69a].

X [GK69]. X-ray [GK69].

References

Anderson:1968:CCC

[AF68] Roger E. Anderson and J. W. Frazer. Computer control in chemistry at the Lawrence Radiation Laboratory. Journal of Computational Physics, 2(4):484–493, June 1968. CODEN JCTPAH.

ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916890051X.

Amsden:1968:TTN

[AH68] Anthony A. Amsden and Francis H. Harlow. Transport of turbulence in numerical fluid dynamics. Journal of Computational Physics, 3(1):94–110, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900089.

Aharoni:1969:C

[Aha69] Amikam Aharoni. Computation of K_{p}(x). Journal of Computational Physics, 4(2):270–271, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900722.

Alterman:1969:MCF

[AI69] Z. S. Alterman and E. Isaacson. A method for calculating frontal motion I. One dimensional tests. Journal of Computational Physics, 4(1):67–96, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900412.

Abraham:1969:FPC

[AJ69] Farid F. Abraham and Stanley K. Jordan. Fortran program for computing the transient and steady-state modifications of a raindrop size distribution due to evaporation and coalescence. Journal of Computational Physics, 4(1):144–145, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900461.

Anonymous:1967:AIV

[Ano67a] Anonymous. Author index of volume 1. Journal of Computational Physics, 1(4):545, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900605.

Anonymous:1967:E

[Ano67b] Anonymous. Erratum. Journal of Computational Physics, 1(4): 544, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900599.

Anonymous:1968:AIV

[Ano68] Anonymous. Author index of volume 2. Journal of Computational Physics, 2(4):494, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900521.

Anonymous:1969:AIVa

[Ano69a] Anonymous. Author index of volume 3. Journal of Computational Physics, 3(4):547–548, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900631.

Anonymous:1969:AIVb

[Ano69b] Anonymous. Author index of volume 4. Journal of Computational Physics, 4(4):601–602, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900242.

Anderson:1968:CFE

[APR68] James L. Anderson, Stanley Preiser, and Ephraim L. Rubin. Conservation form of the equations of hydrodynamics in curvilinear coordinate systems. Journal of Computational Physics, 2(3):279– 287, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900582.

Anderson:1966:NIS

[APS66] J. L. Anderson, J. K. Percus, and J. K. Steadman. Numerical investigations of a simple model of a one-dimensional fluid. Journal of Computational Physics, 1(1):68–86, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999166900131.

Arakawa:1966:CDL

[Ara66] Akio Arakawa. Computational design for long-term numerical integration of the equations of fluid motion: Two-dimensional incompressible flow. Part I. Journal of Computational Physics, 1 (1):119–143, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999166900155.

Allen:1969:SCC

[AWS69] R. C. Allen, Jr., G. M. Wing, and M. Scott. Solution of a certain class of nonlinear two-point boundary value problems. Journal of Computational Physics, 4(2):250–257, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900709.

Baker:1969:CUP

[Bak69] George A. Baker, Jr. A certain unfolding problem. Journal of Computational Physics, 3(4):486–493, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900576.

Bell:1969:DLS

[BD69] D. H. Bell and L. M. Delves. Discrete least squares method for the solution of the Schrödinger equation: Application to the nuclear three-body problem. Journal of Computational Physics, 3 (4):453–472, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900552.

Bertman:1968:APC

[Ber68] B. Bertman. An Algol program for calibrating resistance thermometers at low temperatures. Journal of Computational Physics, 3(1):157–158, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900120.

Birdsall:1969:CCC

[BF69] Charles K. Birdsall and Dieter Fuss. Clouds-in-clouds, clouds-in-cells physics for many-body plasma simulation. Journal of Computational Physics, 3(4):494–511, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900588.

Boris:1969:DSS

[BG69a] Jay P. Boris and John M. Greene. Determination of subdominant solutions using a partial Wronskian. Journal of Computational Physics, 4(1):30–42, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900382.

Brandt:1969:AAH

[BG69b] A. Brandt and J. Gillis. Asymptotic approach to Hartmann– Poiseuille flows. Journal of Computational Physics, 3(4):523– 538, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900606.

Baer:1967:GCF

[BK67] F. Baer and R. L. King. A general computational form for a class of nonlinear systems incorporating both spectral and finitedifference approximations. Journal of Computational Physics, 2 (1):32–60, August 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900393.

Bellman:1966:IIR

[BKK66] Richard Bellman, Harriet Kagiwada, and Robert Kalaba. Invariant imbedding and radiative transfer in spherical shells. Journal of Computational Physics, 1(2):245–256, November 1966. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999166900052.

Bellan:1967:NRE

[BKKU67] R. Bellan, H. Kagiwada, R. Kalaba, and S. Ueno. Numerical results for the estimation of source distributions from external radiation-field measurements. Journal of Computational Physics, 1(4):457–470, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900526.

Byckling:1969:MCM

[BKKV69] E. Byckling, M. Kaartinen, K. Kajantie, and H. Villanen. A Monte Carlo method for generating peripheral events. Journal of Computational Physics, 4(4):521–530, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900175.

Blatt:1967:PPC

[Bla67] John M. Blatt. Practical points concerning the solution of the Schrödinger equation. Journal of Computational Physics, 1(3): 382–396, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900460.

Boris:1969:OPC

[BR69] J. P. Boris and K. V. Roberts. The optimization of particle calculations in 2 and 3 dimensions. Journal of Computational Physics, 4 (4):552–571, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900205.

Bryan:1969:NMS

[Bry69] Kirk Bryan. A numerical method for the study of the circulation of the world ocean. Journal of Computational Physics, 4 (3):347–376, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900047.

Brubaker:1968:TDD

[BS68] T. A. Brubaker and D. R. Stevens. Time-domain design of finitememory digital filters. Journal of Computational Physics, 2(4): 465–483, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900508.

Boynton:1969:NCS

[BT69] Frederick P. Boynton and Alex Thomson. Numerical computation of steady, supersonic, two-dimensional gas flow in natural coordinates. Journal of Computational Physics, 3(3):379–398, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900771.

Buneman:1967:TRD

[Bun67] O. Buneman. Time-reversible difference procedures. Journal of Computational Physics, 1(4):517–535, June 1967. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999167900563.

Burstein:1966:FDC

[Bur66] Samuel Z. Burstein. Finite-difference calculations for hydrodynamic flows containing discontinuities. Journal of Computational Physics, 1(2):198–222, November 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999166900039.

Byers:1967:NCM

[Bye67] J. A. Byers. A nonlinear computer model applicable to low-beta plasma interchange modes in two dimensions. Journal of Computational Physics, 1(4):496–516, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900551.

Cameron:1966:AEC

[Cam66] I. G. Cameron. An analysis of the errors caused by using artificial viscosity terms to represent steady-state shock waves. Journal of Computational Physics, 1(1):1–20, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916690009X.

Campbell:1968:NSP

[Cam68] J. A. Campbell. Note a short program for Dirac matrix algebra. Journal of Computational Physics, 2(4):412–413, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900466.

Carter:1969:SME

[Car69a] Benjamin P. Carter. S-state matrix elements for muonic molecules. Journal of Computational Physics, 4(1):54–66, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900400.

Carter:1969:MMC

[Car69b] Benjamin P. Carter. Muonic-molecule calculations by variation of functions. Journal of Computational Physics, 4(3):325–346, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900035.

Clementi:1966:ESL

[CD66] E. Clementi and D. R. Davis. Electronic structure of large molecular systems. Journal of Computational Physics, 1(2):223–244, November 1966. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999166900040.

Chen:1969:CIA

[Che69] Reuven Chen. On the computation of the integral appearing in glow curve theory. Journal of Computational Physics, 4(3):415– 418, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900084.

Chorin:1967:NMS

[Cho67] Alexandre Joel Chorin. A numerical method for solving incompressible viscous flow problems. Journal of Computational Physics, 2(1):12–26, August 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/002199916790037X.

Concus:1967:NSN

[Con67] Paul Concus. Numerical solution of the nonlinear magnetostaticfield equation in two dimensions. Journal of Computational Physics, 1(3):330–342, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900435.

Costantino:1969:TDW

[Cos69] Carl J. Costantino. Two dimensional wave propagation through nonlinear media. Journal of Computational Physics, 4(2):147– 170, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900643.

Crowley:1967:PSL

[Cro67a] Barbara K. Crowley. PUFL, an “almost–Lagrangian” gasdynamic calculation for pipe flows with mass entrainment. Journal of Computational Physics, 2(1):61–86, August 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916790040X.

Crowley:1967:SON

[Cro67b] W. P. Crowley. Second-order numerical advection. Journal of Computational Physics, 1(4):471–484, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900538.

Crowley:1968:GNO

[Cro68] W. P. Crowley. A global numerical ocean model: Part I. Journal of Computational Physics, 3(1):111–147, August 1968. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999168900090.

Costello:1967:NSH

[CS67] Frederick A. Costello and George L. Schrenk. Numerical solution to the heat-transfer equations with combined conduction and radiation. Journal of Computational Physics, 1(4):541– 543, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900587.

Daly:1969:TIS

[Dal69] Bart J. Daly. A technique for including surface tension effects in hydrodynamic calculations. Journal of Computational Physics, 4(1): 97–117, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900424.

Debar:1967:DEL

[Deb67] Roger B. Debar. Difference equations for the Legendre polynomial representation of the transport equation. Journal of Computational Physics, 2(2):197–205, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900344.

Delves:1968:REV

[Del68] L. M. Delves. Round-off errors in variational calculations. Journal of Computational Physics, 3(1):17–28, August 1968. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999168900028.

Dorny:1968:FDA

[Dor68] C. Nelson Dorny. Finite-difference approximation of the exterior problem for Poisson’s equation. Journal of Computational Physics, 2(4):363–380, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900430.

Dickmann:1967:FPC

[DS67] D. B. Dickmann and G. E. Schacher. A Fortran program for calculating internal fields in general dipole lattices. Journal of Computational Physics, 2(2):87–89, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900277.

Emery:1968:ESD

[Eme68] Ashley F. Emery. An evaluation of several differencing methods for inviscid fluid flow problems. Journal of Computational Physics, 2 (3):306–331, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900600.

Faccioli:1968:DEM

[FA68] E. Faccioli and A. H.-S. Ang. A discrete Eulerian model of spherically symmetric compressible media. Journal of Computational Physics, 3(2):226–258, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916890020X.

Fosdick:1968:ACW

[FJ68] Lloyd D. Fosdick and Harry F. Jordan. Approximation of a conditional Wiener integral. Journal of Computational Physics, 3(1): 1–16, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900016.

Fromm:1968:MRD

[Fro68] Jacob E. Fromm. A method for reducing dispersion in convective difference schemes. Journal of Computational Physics, 3(2):176– 189, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900156.

Friedl:1967:ULD

[FSD^{+}67] Paul Friedl, Charles Sederholm, William Dye, Hall Crannell, and M. R. Yearian. The use of an on-line digital computer in closedloop high-energy physics. Journal of Computational Physics, 1 (3):305–321, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900411.

Gabriel:1968:CCH

[Gab68] J. R. Gabriel. The calculation of cubic harmonics. Journal of Computational Physics, 2(3):336–338, February 1968. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999168900636.

Gary:1969:CTD

[Gar69] John M. Gary. A comparison of two difference schemes used for numerical weather prediction. Journal of Computational Physics, 4(3):279–305, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900011.

Grant:1966:DRR

[GB66] I. P. Grant and V. M. Burke. The diffuse reflection of radiation by model planetary atmospheres. Journal of Computational Physics, 1 (2):277–302, November 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999166900076.

Gold:1968:EEI

[GB68] Raymond Gold and Edgar F. Bennett. Error estimates for iterative unfolding. Journal of Computational Physics, 3(2):167– 175, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900144.

Gold:1967:MNE

[GCO67] Raymond Gold, Charles E. Cohn, and Ingeborg Olson. Methods for the numerical evaluation of Fourier integrals. Journal of Computational Physics, 1(4):536–540, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900575.

Gerhard:1969:MAP

[GK69] F. Bruce Gerhard, Jr. and J. Lawrence Katz. A method of alphabetic plotting of undistorted Fourier maps in X-ray crystallography. Journal of Computational Physics, 3(3):444–446, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900801.

Gladney:1968:STS

[Gla68] H. M. Gladney. A simple time-sharing monitor system for laboratory automation. Journal of Computational Physics, 2(3):255– 273, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900569.

Gwinn:1968:LCD

[GLSM68] William D. Gwinn, Alan C. Luntz, C. H. Sederholm, and Richard Millikan. On-line control, data collection, and reduction for chemical experiments at Berkeley. Journal of Computational Physics, 2(4):439–464, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900491.

Gentry:1966:EDM

[GMD66] Richard A. Gentry, Robert E. Martin, and Bart J. Daly. An Eulerian differencing method for unsteady compressible flow problems. Journal of Computational Physics, 1(1):87–118, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999166900143.

Grant:1968:NAD

[Gra68] I. P. Grant. Numerical analysis of discrete ordinate methods. Journal of Computational Physics, 2(4):381–402, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900442.

Goldberg:1967:ISEa

[GS67a] Abraham Goldberg and Judah L. Schwartz. Integration of the Schrödinger equation in imaginary time. Journal of Computational Physics, 1(3):433–447, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900496.

Goldberg:1967:ISEb

[GS67b] Abraham Goldberg and Judah L. Schwartz. Integration of the Schrödinger equation in imaginary time. II. Journal of Computational Physics, 1(3):448–453, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900502.

Grimm:1969:NMN

[GS69] R. Grimm and R. G. Storer. A new method for the numerical solution of the Schrödinger equation. Journal of Computational Physics, 4(2):230–249, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900692.

Harlow:1968:NCA

[HA68] Francis H. Harlow and Anthony A. Amsden. Numerical calculation of almost incompressible flow. Journal of Computational Physics, 3(1):80–93, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900077.

Hagin:1968:CES

[Hag68] F. G. Hagin. Computation of eigenvalues for second-order differential equations using imbedding techniques. Journal of Computational Physics, 3(1):46–57, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900053.

Hamilton:1968:CCC

[Ham68] Walter C. Hamilton. Computers in chemistry conference: Computers, chemists, and crystallography. Journal of Computational Physics, 2(4):417–438, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916890048X.

Harrison:1969:CMS

[Har69] George Harvey Harrison. A compact method for symbolic computation of the Riemann tensor. Journal of Computational Physics, 4 (4):594–600, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900230.

Hategan:1969:RSC

[Hat69] C. Hategan. Resolution of a system of coupled Schrödinger equations. Journal of Computational Physics, 4(1):43–53, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900394.

Hirt:1967:GCP

[HH67] C. W. Hirt and Francis H. Harlow. A general corrective procedure for the numerical solution of initial-value problems. Journal of Computational Physics, 2(2):114–119, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900290.

Hohl:1969:CMD

[HH69] Frank Hohl and R. W. Hockney. A computer model of disks of stars. Journal of Computational Physics, 4(3):306–324, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900023.

Hirt:1968:HST

[Hir68] C. W. Hirt. Heuristic stability theory for finite-difference equations. Journal of Computational Physics, 2(4):339–355, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900417.

Houghton:1969:NML

[HJ69] David D. Houghton and Walter L. Jones. A numerical model for linearized gravity and acoustic waves. Journal of Computational Physics, 3(3):339–357, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900758.

Harris:1969:IEA

[HM69] Frank E. Harris and H. H. Michels. Integrals for electron-atom scattering calculations. Journal of Computational Physics, 4(4): 579–593, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900229.

Higbie:1968:VPO

[HP68] Leslie C. Higbie and Myron N. Plooster. Variable pseudoviscosity in one-dimensional hyperbolic difference schemes. Journal of Computational Physics, 3(1):154–156, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900119.

Heldmann:1968:CAC

[HS68a] G. Heldmann and P. Schnupp. Computer-aided construction of symmetry adapted wave functions and matrix elements. Journal of Computational Physics, 3(2):208–216, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900181.

Hicks:1968:AMC

[HS68b] Bruce L. Hicks and Margaret A. Smith. On the accuracy of Monte Carlo solutions of the nonlinear Boltzmann Equation. Journal of Computational Physics, 3(1):58–79, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900065.

Hirt:1968:FSS

[HS68c] C. W. Hirt and J. P. Shannon. Free-surface stress conditions for incompressible-flow calculations. Journal of Computational Physics, 2(4):403–411, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900454.

Innanen:1966:EPI

[Inn66] Kimmo A. Innanen. An example of precise interpolation with a spline function. Journal of Computational Physics, 1(2):303– 304, November 1966. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999166900088.

Joshi:1969:RAE

[JK69] Bhairav D. Joshi and K. L. Kapoor. The recursive and analytical evaluation of atomic correlation integrals. Journal of Computational Physics, 3(3):447–450, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900813.

Jackson:1968:TCO

[JPW68] E. Atlee Jackson, John R. Pasta, and John F. Waters. Thermal conductivity of one-dimensional lattices. Journal of Computational Physics, 2(3):207–227, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900533.

Kalos:1966:SWF

[Kal66] M. H. Kalos. Stochastic wave function for atomic helium. Journal of Computational Physics, 1(2):257–276, November 1966. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999166900064.

Krell:1968:BSS

[KE68] M. Krell and T. E. O. Ericson. The bound-state solution of wave equations for real or complex eigenvalues. Journal of Computational Physics, 3(2):202–207, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916890017X.

Kelman:1967:PCE

[Kel67] Robert B. Kelman. The potential in a cylindrical electrode partially submerged in a perfectly conducting medium. Journal of Computational Physics, 2(2):120–125, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900307.

Killeen:1968:NSF

[KF68] John Killeen and Archer H. Futch. Numerical solution of the Fokker–Planck equations for a hydrogen plasma formed by neutral injection. Journal of Computational Physics, 2(3):236–254, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900557.

Kasahara:1969:END

[KH69] Akira Kasahara and David D. Houghton. An example of nonunique, discontinuous solutions in fluid dynamics. Journal of Computational Physics, 4(3):377–388, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900059.

Kagiwada:1967:IVM

[KK67] Harriet Kagiwada and R. E. Kalaba. Initial-value methods for the basic boundary-value problem and integral equation of radiative transfer. Journal of Computational Physics, 1(3):322– 329, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900423.

Kagiwada:1968:FEA

[KKS68] H. Kagiwada, R. Kalaba, and R. Segerblom. Flux equivalences among Rayleigh, isotropic, and other scattering models. Journal of Computational Physics, 3(2):159–166, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900132.

Kusic:1969:SDM

[KL69] G. L. Kusic and A. Lavi. Stability of difference methods for initial-value type partial differential equations. Journal of Computational Physics, 3(3):358–378, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916990076X.

Kent:1969:NSV

[KM69a] Gordon Kent and Joseph Mautz. The numerical solution of a Volterra integral equation. Journal of Computational Physics, 3 (3):399–415, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900783.

Kopp:1969:VPC

[KM69b] Marcel Kopp and John H. Mackey. A versatile program for computing paramagnetic resonance spectra from a parametric spin-Hamiltonian. Journal of Computational Physics, 3(4):539– 543, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900618.

Kopp:1969:CAP

[Kop69] Marcel Kopp. Computation and analysis of pulsed-NMR line shapes of crystals. Journal of Computational Physics, 4(2):195– 203, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900667.

Killeen:1966:CSF

[KR66] John Killeen and Shirley L. Rompel. A computation for studying the formation of the relativistic electron layer in astron. Journal of Computational Physics, 1(1):29–50, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999166900118.

Kusic:1969:SNM

[Kus69] George Kusic. On stability of numerical methods for systems of initial-value partial differential equations. Journal of Computational Physics, 4(2):272–275, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900734.

Lapidus:1967:DSC

[Lap67] Arnold Lapidus. A detached shock calculation by second-order finite differences. Journal of Computational Physics, 2(2):154–177, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900320.

Lathrop:1969:SDT

[Lat69] K. D. Lathrop. Spatial differencing of the transport equation: Positivity vs. accuracy. Journal of Computational Physics, 4(4): 475–498, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900151.

Laura:1969:C

[Lau69] Patricio A. Laura. Comment. Journal of Computational Physics, 4(2):276–277, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900746.

Lathrop:1966:NSB

[LC66] K. Lathrop and B. Carlson. Numerical solution of the Boltzmann Transport Equation. Journal of Computational Physics, 1 (2):173–197, November 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999166900027.

Libby:1968:AQL

[LC68] Paul A. Libby and Karl K. Chen. Application of quasilinearization to an eigenvalue problem arising in boundary-layer theory. Journal of Computational Physics, 2(4):356–362, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900429.

Larcher:1969:DRS

[LC69] J. F. Larcher and D. P. Chong. On degenerate Rayleigh–Schrödinger perturbation. Journal of Computational Physics, 4(3):411–414, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900072.

Lester:1968:VMN

[Les68] William A. Lester, Jr. De Vogelaere’s method for the numerical integration of second-order differential equations without explicit first derivatives: Application to coupled equations arising from the Schrödinger equation. Journal of Computational Physics, 3 (2):322–326, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900259.

Levine:1967:DMT

[Lev67] M. J. Levine. Dirac matrix and tensor algebra on a digital computer. Journal of Computational Physics, 1(3):454–455, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900514.

Lick:1969:DTN

[Lic69] Dale W. Lick. A divergence theorem for a nonlinear Dirichlet problem. Journal of Computational Physics, 4(1):142–143, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002199916990045X.

Lawande:1969:MCE

[LJS69a] S. V. Lawande, C. A. Jensen, and H. L. Sahlin. Monte Carlo evaluation of Feynman Path Integrals in imaginary time and spherical polar coordinates. Journal of Computational Physics, 4(4): 451–464, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900138.

Lawande:1969:MCI

[LJS69b] S. V. Lawande, C. A. Jensen, and H. L. Sahlin. Monte Carlo integration of the Feynman propagator in imaginary time. Journal of Computational Physics, 3(3):416–443, February 1969. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999169900795.

Lyons:1969:EIR

[LN69] J. D. Lyons and R. K. Nesbet. Evaluation of integrals required in scattering theory: I. Non-exchange type integrals. Journal of Computational Physics, 4(4):499–520, December 1969. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999169900163.

Matin:1968:NIE

[Mat68] Shaikh A. Matin. Numerical iteration of an elliptic mixed boundaryvalue problem in a region with curved boundary. Journal of Computational Physics, 3(2):327–330, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900260.

Matthes:1969:MCP

[Mat69] W. Matthes. Model calculation of phase transitions with Monte Carlo. Journal of Computational Physics, 4(4):431–450, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900126.

Minicozzi:1969:DIE

[MB69] William P. Minicozzi and Dan F. Bradley. On the determination of interaction energy surfaces I. Formic and acetic acid dimers. Journal of Computational Physics, 4(1):118–137, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900436.

Miller:1967:EMT

[Mil67] Richard H. Miller. An experimental method for testing numerical stability in initial-value problems. Journal of Computational Physics, 2(1):1–7, August 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900356.

Mjolsness:1968:MSH

[MR68] R. C. Mjolsness and H. M. Ruppel. Multiple solutions of Hartree– Fock equations. Journal of Computational Physics, 3(2):259– 272, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900211.

Ng:1968:DSS

[Ng68] Edward W. Ng. On the direct summation of series involving higher transcendental functions. Journal of Computational Physics, 3 (2):334–338, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900296.

Nagaraja:1969:ALW

[NH69] K. S. Nagaraja and James P. Hudson. On the application of Luke’s weight coefficients for evaluating the linearized Rayleigh problem in a rarefied gas flow. Journal of Computational Physics, 3(3): 451–452, February 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900825.

OBrien:1969:DNS

[O’B69] James J. O’Brien. A direct numerical solution to a one-dimensional Helmholtz equation. Journal of Computational Physics, 3(4):544– 546, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199916990062X. See comments [Yee70].

Petschek:1968:DET

[PH68] A. G. Petschek and M. E. Hanson. Difference equations for twodimensional elastic flow. Journal of Computational Physics, 3 (2):307–321, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900247.

Pierce:1967:MAI

[Pie67] Allan D. Pierce. The multilayer approximation for infrasonic wave propagation in a temperature- and wind-stratified atmosphere. Journal of Computational Physics, 1(3):343–366, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999167900447.

Pan:1967:EME

[PR67] Yuh-Kang Pan and Max T. Rogers. The evaluation of matrix elements of the high-resolution NMR spin-coupling Hamiltonian using a digital computer. Journal of Computational Physics, 2(1): 27–31, August 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900381.

Rubin:1967:DMI

[RB67] Ephraim L. Rubin and Samuel Z. Burstein. Difference methods for the inviscid and viscous equations of a compressible gas. Journal of Computational Physics, 2(2):178–196, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999167900332.

Rogers:1969:NSW

[RB69] E. H. Rogers and D. W. Beard. A numerical study of widegap Taylor vortices. Journal of Computational Physics, 4(1):1– 18, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900369.

Rigler:1969:CSV

[Rig69] A. K. Rigler. A choice of starting vectors in relaxation methods. Journal of Computational Physics, 4(3):419–423, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900096.

Reeves:1969:SSE

[RO69] M. Reeves III and L. W. Owen. Solution of a Schrödinger equation with a nonlocal potential. Journal of Computational Physics, 4(4):572–578, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900217.

Rajaiah:1968:EBC

[RR68] K. Rajaiah and Akella Kameswara Rao. Effect of boundary condition description on convergence of solution in a boundaryvalue problem. Journal of Computational Physics, 3(2):190–201, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900168.

Rusbridge:1968:RRS

[Rus68] M. G. Rusbridge. On the realization of random surfaces. Journal of Computational Physics, 2(3):288–305, February 1968. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999168900594.

Sauerwein:1967:NCM

[Sau67] Harry Sauerwein. Numerical calculations of multidimensional and unsteady flows by the method of characteristics. Journal of Computational Physics, 1(3):406–432, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900484.

Schwartz:1966:ICB

[Sch66] Charles Schwartz. Information content of the Born series. Journal of Computational Physics, 1(1):21–28, August 1966. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999166900106.

Schwartz:1967:NTM

[Sch67] Charles Schwartz. Numerical techniques in matrix mechanics. Journal of Computational Physics, 2(2):90–113, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999167900289.

Schwartz:1969:NIA

[Sch69a] Charles Schwartz. Numerical integration of analytic functions. Journal of Computational Physics, 4(1):19–29, June 1969. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0021999169900370.

Schwartz:1969:VPI

[Sch69b] Charles Schwartz. Variational principles for integrals. Journal of Computational Physics, 3(4):512–522, April 1969. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002199916990059X.

Suarez:1969:FPA

[SG69] Carlos B. Suárez and Mario Gallardo. A Fortran program to analyze the rotational structure of diatomic molecules. Journal of Computational Physics, 4(3):424–426, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900102.

Schaefer:1968:CUA

[SH68] Henry F. Schaefer III and Frank E. Harris. Construction and use of atomic L-S eigenfunctions. Journal of Computational Physics, 3(2):217–225, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900193.

Shaw:1967:NSP

[Sha67] Tzu-Sien Shaw. Numerical solution of plane viscous shock reflections. Journal of Computational Physics, 1(3):367–381, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999167900459.

Sloan:1968:NEN

[Slo68a] I. H. Sloan. Note errors in the Numerov and Runge–Kutta methods. Journal of Computational Physics, 2(4):414–416, June 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900478.

Sloan:1968:NEP

[Slo68b] I. H. Sloan. The numerical evaluation of principal-value integrals. Journal of Computational Physics, 3(2):332–333, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900284.

Sloan:1968:MNS

[Slo68c] Ian H. Sloan. Method for the numerical solution of linear secondorder differential equations. Journal of Computational Physics, 3(1):40–45, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900041.

Southwell:1969:FED

[Sou69] W. H. Southwell. Fitting experimental data. Journal of Computational Physics, 4(4):465–474, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916990014X.

Sankar-Rao:1968:MSS

[SR68] M. Sankar-Rao. A method of solution for a system of two secondorder ordinary differential equations arising in the theory of the mean atmospheric waves. Journal of Computational Physics, 2 (3):228–235, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900545.

Stein:1967:CPA

[Ste67] Josef Stein. Computational problems associated with Racah algebra. Journal of Computational Physics, 1(3):397–405, February 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999167900472.

Steele:1968:NCE

[Ste68] Charles W. Steele. Numerical computation of electric and magnetic fields in a uniform waveguide of arbitrary cross section. Journal of Computational Physics, 3(1):148–153, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999168900107.

Swanson:1969:CAD

[Swa69] Stanley M. Swanson. Computer algorithms for Dirac algebra. Journal of Computational Physics, 4(2):171–194, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900655.

Taylor:1968:CSP

[Tay68] David B. Taylor. The calculation of steady plane supersonic gas flows containing an arbitrarily large number of shocks. Journal of Computational Physics, 3(2):273–290, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900223.

Taylor:1969:TTD

[Tay69] P. J. Taylor. A technique for treating Dirichlet conditions at infinity in numerical field problems. Journal of Computational Physics, 4(1):138–141, June 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900448.

Textor:1969:SCV

[TLF69] R. E. Textor, D. W. Lick, and G. J. Farris. Solution of confined vortex problems. Journal of Computational Physics, 4(2): 258–269, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900710.

Trower:1966:FPC

[Tro66a] W. Peter Trower. A FORTRAN program for calculating kinematic and dynamic quantities of particle interactions and decays. Journal of Computational Physics, 1(1):144–145, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999166900167.

Trower:1966:FSC

[Tro66b] W. Peter Trower. A FORTRAN subroutine for calculating the range-energy relation of charged particles in chemical elements. Journal of Computational Physics, 1(1):145–147, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999166900179.

Talley:1969:MCA

[TW69] Wilson K. Talley and Stephen Whitaker. Monte Carlo analysis of Knudsen flow. Journal of Computational Physics, 4(3):389– 410, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900060.

Usher:1967:CSI

[Ush67] Peter D. Usher. Coordinate stretching and interface location I. Star models in quasi-static equilibrium. Journal of Computational Physics, 2(1):8–11, August 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999167900368.

Usher:1968:CSI

[Ush68] Peter D. Usher. Coordinate stretching and interface location II. A new PL expansion. Journal of Computational Physics, 3(1): 29–39, August 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002199916890003X.

vandeHulst:1968:AFM

[vdH68] H. C. van de Hulst. Asymptotic fitting, a method for solving anisotropic transfer problems in thick layers. Journal of Computational Physics, 3(2):291–306, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900235.

Verdier:1969:SMS

[Ver69] Peter H. Verdier. A simulation model for the study of the motion of random-coil polymer chains. Journal of Computational Physics, 4(2):204–210, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900679.

Viecelli:1969:MIA

[Vie69] J. A. Viecelli. A method for including arbitrary external boundaries in the MAC incompressible fluid computing technique. Journal of Computational Physics, 4(4):543–551, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900199.

vanLeer:1969:SDS

[vL69] B. van Leer. Stabilization of difference schemes for the equations of inviscid compressible flow by artificial diffusion. Journal of Computational Physics, 3(4):473–485, April 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999169900564.

Wayland:1968:PCP

[Way68] J. R. Wayland. A program to calculate particle spectra produced in high-energy proton-proton collisions by the two-temperature statistical model. Journal of Computational Physics, 3(2):331, October 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900272.

Williams:1968:FPC

[WCT68] I. R. Williams, M. Craig, Jr., and C. L. Thompson. A Fortran program for calculating the solid angle subtended by one circular disk at another. Journal of Computational Physics, 2(3):332– 333, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999168900612.

Wendroff:1969:DSR

[Wen69] Burton Wendroff. A difference scheme for radiative transfer. Journal of Computational Physics, 4(2):211–229, August 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0021999169900680.

Williamson:1966:SDA

[Wil66] David Williamson. Stability of difference approximations to certain partial differential equations of fluid dynamics. Journal of Computational Physics, 1(1):51–67, August 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916690012X.

Winslow:1966:NSQ

[Win66] Alan M. Winslow. Numerical solution of the quasilinear Poisson equation in a nonuniform triangle mesh. Journal of Computational Physics, 1(2):149–172, November 1966. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999166900015.

Wind:1968:PMF

[Win68] H. Wind. Processing magnetic field data. Journal of Computational Physics, 2(3):274–278, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999168900570.

Wolf:1968:SFS

[Wol68] Kurt Bernardo Wolf. A set of Fortran subroutines for handling bases of group representations. Journal of Computational Physics, 2(3):334–335, February 1968. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999168900624.

Yanowitch:1969:NSV

[Yan69] M. Yanowitch. A numerical study of vertically propagating waves in a viscous isothermal atmosphere. Journal of Computational Physics, 4(4):531–542, December 1969. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999169900187.

Yee:1970:CDN

[Yee70] Samuel Y. K. Yee. Comments on “A direct numerical solution to a one-dimensional Helmholtz equation”. Journal of Computational Physics, 6(2):336–337, October 1970. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0021999170900331.

See [O’B69].

Yuta:1967:DMA

[Yut67] H. Yuta. Determination of momentum and angles of bubblechamber tracks in a nonuniform magnetic field. Journal of Computational Physics, 1(4):485–495, June 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002199916790054X.

Zaker:1969:CCE

[Zak69] T. A. Zaker. Calculation of the complementary error function of complex argument. Journal of Computational Physics, 4(3):427– 430, October 1969. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0021999169900114.

Zabusky:1967:DNL

[ZD67] Norman J. Zabusky and Gary S. Deem. Dynamics of nonlinear lattices I. Localized optical excitations, acoustic radiation, and strong nonlinear behavior. Journal of Computational Physics, 2 (2):126–153, November 1967. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.sciencedirect.

com/science/article/pii/0021999167900319.